OpenStudy (anonymous):

simplify the alabraic equation: 2(5x+4)+3x-(7-x)

OpenStudy (m):

you want to group all numbers with x

OpenStudy (amistre64):

distribute thru the paranthesis and then combine like terms

OpenStudy (m):

^^

OpenStudy (anonymous):

ok seriously i suck at math that makes no sense to me i have no idea how to do that

OpenStudy (amistre64):

lets figure out what distribute means then :)

OpenStudy (amistre64):

would you agree that (5+3) = 8?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then by all means: 4(5+3) = 32 right? becasue it is simply 4(8)

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

to distribute means to take that 4 and distribute it thru to the 5 and 3 like this: 4(5) + 4(3) = 32 20 + 12 = 32 32 = 32 ...did you see that?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

4(x+3) would distribute like this: 4(x) + 4(3) 4x + 21 right?

OpenStudy (anonymous):

sure

OpenStudy (amistre64):

well, 4(3) = 12, not 21..typoed it

OpenStudy (amistre64):

can you distribute this for me then? 2(5x+4)

OpenStudy (anonymous):

2 and the 5 is = 10 4 and 5 is= 20 so is it 30

OpenStudy (amistre64):

good job!! you were close; but you forgot to include the 'x' part of it :) 2(5x+4) 2(5x) + 2(4) 10x + 8

OpenStudy (amistre64):

a little more practice and you should have it down pat

OpenStudy (anonymous):

oh so its 18 im gettng it

OpenStudy (amistre64):

not quite 18; the x represent something that we know nothing about yet; suppose it represent oranges would 10 oranges and 8 cents = 18?? or would it simply be: 10 oranges and 18 cents?

OpenStudy (amistre64):

typoed again...

OpenStudy (amistre64):

not quite 18; the x represent something that we know nothing about yet; suppose it represent oranges would 10 oranges and 8 cents = 18?? or would it simply be: 10 oranges and 18 cents?

OpenStudy (amistre64):

.....computer spazed oon me lol

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

10x is just 10x; theres nothing more to do with it. 10x + 8 is as far as we can go with it

OpenStudy (amistre64):

try to distribute this one for me: -1(7-x)

OpenStudy (anonymous):

ok slowly but surely getting it

OpenStudy (anonymous):

-1+7x

OpenStudy (amistre64):

still got some ironing out to do; but the concepts are there :)

OpenStudy (amistre64):

try to rewrite it like this: -1(7) + -1(-x)

OpenStudy (amistre64):

-7 + x right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

10x+8 +3x -7 +x is what we have come to in your question then; now we add up all the x's to get what? 10x +3x +x = how many x's?

OpenStudy (anonymous):

3

OpenStudy (amistre64):

3x means: x+x+x x mean: x 10x means: x+x+x+x+x+x+x+x+x+x when we add up all these x's how many are there? (x) + (x+x+x) + (x+x+x+x+x+x+x+x+x+x) ?

OpenStudy (anonymous):

14

OpenStudy (amistre64):

thats correct :) 14x is the combined number of x's :)

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

14x +(8-7) =? what is 8-7?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

then we have come to our answer and cant go any further: 14x + 1 is the answer :)

OpenStudy (anonymous):

oh i see

OpenStudy (amistre64):

practice at it; the hardest part really is getting the distributive stuff down pat.

OpenStudy (anonymous):

uh huh i will